OpenStudy (anonymous):

The table below represents a linear function f(x) and the equation represents a function g(x): x f(x) −1 −9 0 −1 1 7 g(x) = 3x − 2 Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). (6 points) Part B: Which function has a greater y-intercept? Justify your answer. (4 points)

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

computer the slope of f(x)

OpenStudy (anonymous):

@jenny1994 computer?

OpenStudy (anonymous):

compute*

OpenStudy (anonymous):

@jenny1994 i don't know how to find the slope without the graph

OpenStudy (anonymous):

\(m=\dfrac{y_2-y_1}{x_2-x_1}\) take any 2 points

OpenStudy (anonymous):

so the first two on the table

OpenStudy (anonymous):

so we have \((x_1,y_1)=(−1 −9)\\ (x_2,y_2) =(0, −1)\) plug and play

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

8 You can also look at the table. y goes up by 8 units with every 1 x unit if you want the work,